# Learning and Transferring Physical Models through Derivatives

### v2

@@ 4.2 Generalization to new ODE/PDE parameters
We consider the dynamical system of a damped pendulum with state equation (DP) in table 1. Additional details on the experimental setup are available in Section D.4. We sampled a total of $50$ trajectories from different starting conditions $(u_{0},\dot{u}_{0})$, which act as parameters: $30$ are reserved for training, $10$ for validation, and $10$ for testing. Each trajectory was sampled every $\Delta t=0.01s$. The MLPs, which predict $u(t;u_{0},\dot{u}_{0})$ from $(t,u_{0},\dot{u}_{0})$, have 4 MLP with $50$ units each and $\tanh$ activation, and are trained for $200$ epochs with batch size $64$. Initial states are sampled randomly in $(u_{0},\dot{u}_{0})\in[-\frac{\pi}{2},\frac{\pi}{2}]\times[-1.5,1.5]$. In addition to the previously-used metrics and the angular velocities $\|\dot{u}-\hat{\dot{u}}\|_{2}$, we also monitor:
$\circ$ $\|\mathcal{G}[\hat{u}]\|_{2}$ (formally defined in Appendix D.4.1), the PDE residual which describes continuity and differentiability of trajectories with respect to the initial conditions. High values of $\|\mathcal{G}[\hat{u}]\|_{2}$ imply predicted trajectories do not change smoothly when varying $(u_{0},\dot{u}_{0})$, which is not consistent with the real system.
@@
For this task, we adapted Lagrangian Neural Networks (LNN) (Cranmer et al., 2019) and Hamiltonian Neural Networks (HNN) (Greydanus et al., 2019) to the damped pendulum case. We train them on the conservative part of the field (see D.4.2 for details), where they excel as they are specifically designed for conservative fields. Unlike DERL, they also require trajectories to be calculated through external solvers. Remarkably, DERL outperforms both LNN and HNN (figure 4(b)) on the learned field: DERL scores an error of $0.4445$, against $0.4470$ and $0.4496$ for HNN and LNN, respectively.
@@
Inspired by Penwarden et al. (2023), we consider a parametric experiment on the Allen-Cahn equation (AC). The parametric component is the external forcing $f$, calculated from the analytical solution $u(x,y;\xi)=e^{-\xi(x+0.7)}\sin(\pi x)\sin(\pi y)$, with parameter $\xi\in\mathbb{R}$. We reserved $8$ values of $\xi$ for training, $2$ for validation, and $5$ for testing, and we used $1000$ collocation points in the domain. The architecture is the same as in the Allen-Cahn experiment in Section 4.1. For further details and the complete setup, see Appendix D.5. Numerical results on the testing parameters are reported in Table 5, while further results are available in Appendix D.5.1. DERL always scores the best results, highlighting strong generalization abilities to unseen parameterization of the PDE. OUTL underperforms in both metrics, showcasing that learning from the raw data $u$ often leads to misleading solutions that do not follow the underlying physical laws. SOB underperforms as it again struggles to optimize a loss with both $u$ and $\mathrm{D}u$.
@@
Double-parameter Allen-Cahn equation.
We now increase complexity by considering a dual-parameter version of the Allen-Cahn equation (Zou & Karniadakis, 2023). The analytical solution is given by $u(x,y;\xi_{1},\xi_{2})=\sum_{j=1}^{2}\xi_{j}\frac{\sin(j\pi x)\sin(j\pi y)}{j^{2}}$, from which the external forcing $f$ is calculated. We selected $15$ random couples of the parameters $(\xi_{1},\xi_{2})\in[0,1]^{2}$ for training, $5$ for validation and $5$ for test. Numerical results for the test parameters are reported in Table 5. Once again, DERL is consistently the best model for every couple of parameters (see Appendix D.5.2), performing at least two times better in predicting $u$ for unseen ones. Learning $u$ through OUTL leads to very poor performance in terms of physical consistency, even with the additional help of the PINN loss.
@@ 5 Transferring physical knowledge
We have shown that DERL is capable of extracting the relevant physical features from data.We now show how we can take advantage of this to transfer physical information between models through distillation (Hinton et al., 2015), eventually building physical models incrementally. To the best of our knowledge, we are the first to leverage distillation for this purpose.
We set up our transfer experiments by dividing the whole training experience into 2 or 3 incremental steps. Each step extends the physical problem to i) new portions of the time domain, ii) to new portions of the space domain, and iii) to new parameterizations of the PDE.
@@
As a comparison, we train two additional PINN models. PINN-no-distillation updates the teacher with the PINN loss on the new domain, but without distillation on the previous domain, and is expected to forget previous information. PINN-full is a PINN trained on the full domain $\Omega$ without the iterative distillation process. This model is trained for the total sum of epochs of the other methodologies for a fair comparison. In continual learning, this joint training method is often taken as an upper bound for the expected performance of the final continual learning model Masana et al. (2023).
@@ 5.1 Transfer to wider time horizons
Figure 5: Prediction errors for the KdV transfer experiment.
In this first experiment with transferring, we aim to extend the working domain of a PINN in time. We consider the 1D Korteweg-de Vries equation (KdV) in Table 1. The spatial domain is the interval [1,1][-1,1]. while the entire time domain is $[0,1]$. We first train a PINN in the time interval $[0,0.35]$ and we collect its evaluations $\hat{u},\mathrm{D}\hat{u}$ on the training collocation points in a dataset. Then, a second model is trained with the combination of the PINN loss on the second time interval $[0.35,0.7]$, and a distillation loss on the old collocation points. This procedure aims to transfer old knowledge to this new model, while discovering the solution in a new interval of the time domain. Hence, this model should be able to learn the solution in the time-space domain $(t,x)\in[0,0.7]\times[-1,1]$. Finally, we repeat this procedure one more time, extending the model to the time interval $[0.7,1]$. As we will see, the final model has correctly learned the solution in the entire time-space domain of $0,1]\times[-1,1]$. For this task, each model is an MLP with $9$ layers of $50$ units and $\tanh$ activation. Training phases last $30.000,30.000$, and $40.000$ steps each with the Adam optimizer and exponentially decaying lr, starting from $0.001$ with factor $0.9$. The PINN-full and PINN-no-distillation models train for the sum of these steps, that is $100.000$, with the same optimizer and learning rate schedule.
Numerical results are available in Table 6. DERL outperforms all other methodologies in learning the solution to the PDE, either starting from-scratch or by fine-tuning the teacher model. Most importantly, our incremental methodology coupled with the DERL distillation loss, improves the results by three times compared to a standard PINN. Figure 5 shows the prediction error in the whole domain as a function of the cumulative training steps. These losses follow the typical behavior of continual learning, where the loss increases at the beginning of each phase and subsequently decreases, with the difference that performance is measured on the entire final task. The results clearly show that the PINN reaches a minimum at the beginning of training and cannot improve, while our incremental approach performs almost an order of magnitude better in the end. We notice that, even for transferring, it is not enough to distill the model’s outputs, both in terms of consistency and prediction error. Derivatives are both necessary and sufficient to distill the physical knowledge correctly. Finally, we see that incrementally training a PINN without distilling previous knowledge leads to forgetting of the learned solution. See Appendix E.3 for additional figures.
5.2 Transfer to new domains
We leverage our transfer protocol to extend the working domain of a model (Figure 1). We consider the Allen-Cahn equation (AC) from Section 4.1. We first train a PINN in the lower-left quadrant of the domain ($1,0]\times[-1,0]$). We evaluate $\hat{u},\mathrm{D}\hat{u}$ on the training collocation points. Then, a second model is trained with a combination of the PINN loss on the lower-right quadrant ($0,1]\times[-1,0]$), and a distillation loss for the old points. The resulting model should be able to learn the solution in the lower half of the domain ($1,0]\times[1,0]$). The same procedure is repeated to extend the working domain of the model to the upper half ($1,1]\times[0,1]$). The final model should be a good predictor in the entire domain ($1,1]\times[-1,1]$). We used the same architecture as in Section 4.1 and $100$ epochs per step with the BFGS optimizer (Appendix E.4).
@@ 5.3 Transfer to new PDE parameters
Table 6 reports the $L^{2}$ errors and PDE residuals on the test parameters for both the from-scratch and continual cases. DERL proves to be the best distillation protocol. It also surpasses all PINN models, except for the full PINN on the PDE residual metric, whilst being the closest to match its performance. This again shows our method’s potential to transfer physical information across different instances of PDEs.
@@ 5.4 Distillation of higher-order derivatives
Finally, we investigate the effects of distilling higher-order derivatives. We consider the Korteweg-de Vries (KdV) equation (KdV in table 1), a third-order non-linear PDE. A pre-trained PINN teacher distills knowledge to students with identical architecture. We tested Hessian learning (HESL), which distills second-order derivatives and its combination with DERL (DERL+HESL) and SOB (SOB+HESL). The full setup and results can be found in section E.3. Here, we summarize the main findings. We observe that OUTL fails to distill effectively, as its PDE residual of $17.366$ is two orders of magnitude larger than the other models. Interestingly, the best-performing methods, HESL and DERL+HESL, did not access $u$ directly. They achieve the lowest PDE residual ($0.1915$ and $0.1931$, respectively) which is comparable to the PINN PDE residual of $0.1664$. Both approaches score the same error on $u$ as the teacher model, and even a lower BC loss.
@@ 6 Conclusion
We proposed DERL, a supervised methodology to learn physical systems from their partial derivatives. We showed theoretically and experimentally that our method successfully learns the solution to a problem and remains consistent with its physical constraints, outperforming other supervised learning methods. We also found DERL to be effective in transferring physical knowledge across models through distillation. We leveraged this capability to extend the working domain of a PINN and to generalize better across different PDE parameterizations. Finally, we showed the advantage of distilling higher-order derivatives.
Our work shows promising intersections with continual learning, as DERL’s distillation allowed us to build physical models incrementally, showing minimal to no forgetting of previous knowledge. In the future, learning physical and dynamical systems may not always require starting from scratch. Rather, the learning process may be based on the continual composition and integration of physical information across different expert models. A more general and flexible paradigm that could lead to data-efficient and sustainable deep learning solutions.
@@ Appendix A Implementation Details and Setups
In theory, the augmented boundary should contain the values of $u(T,{\bm{x}})$ at the final time $T$. However, PDE problems are usually defined without a final condition, and it is more interesting not to have one to see which methodology can propagate physical information better. For example, SOTA methods as PINN do not include them (Raissi et al., 2019). Therefore, we decided not to include them as well, leading to a harder experimental setup than our theoretical results. For example, the experimental results in Appendix D.2 where the PINN failed to propagate the solution from $t=0$ to higher times, confirm the findings of Wang et al. (2022a) of PINNs being biased during training. However, DERL did not suffer from this issue.
@@ A.1 Computational time
Table 7: Computational time of 11 epoch for each experiment with the Adam optimizer. Allen-Cahn results are for the experimental setup in Appendix D.1. For NCL we reported the computational time of $100$ steps due to the definition and length of one epoch as in Richter-Powell et al. (2022).
@@ A.2 Model tuning
For the pendulum experiments that involve full trajectories, $30$ trajectories are for training and $10$ for validation, while $10$ unseen trajectories are used for model testing. For the pendulum interpolation task in Section D.4, a total of $40$ trajectories are used for in the tuning phase: points with times between $t=3.75$ and $t=6.25$ are unseen by the models and used for tuning/validation, while $10$ unseen complete trajectories are used for testing. For the parametric Allen-Cahn experiment with a single parameter, we calculated the solution on a set of $1000$ collocation points for $13$ values of $\xi\in[0,\pi]$: $8$ are used for training, $2$ for validation and $5$ for testing. Similarly, for the experiment with two parameters, we sampled $25$ couples of $(\xi_{1},\xi_{2})\in[0,1]^{2}$ with a $60/20/20$ split.
@@ Appendix B Proofs of the Theoretical Statements
We show the results for time-independent PDEs, for time-dependent PDEs it is sufficient to consider the extended domain $\tilde{\Omega}=[0,T]\times\Omega$, in which the initial (and final) conditions become boundary conditions. We also show the result for functions with one component. The result can be generalized by applying it to each component.
Let $\hat{u}$ be the trained Neural network. From theorem 3.2, we have that the network converges to the solution of the PDE $\bm{u}$} in $W^{1,2}(\Omega)$ and that the two functions coincide almost everywhere, since $\mathrm{D}(\hat{u}-u)=0$. From this, we deduce that the PDE is satisfied by $\hat{u}$ and the boundary conditions are satisfied as well as per theorem 3.2. ∎
@@
In the case of time-dependent PDEs, we will not provide the final conditions at $t=T$, which are part of the "boundary" of the extended domain and are necessary for theorem 3.3. Since our experiments are on regular functions, this has proven not to be an issue and the neural network still converges.
Appendix C On the use of Empirical Derivatives
As discussed in Section 3, when analytical or true derivatives are not available, we use empirical ones via finite differences (Anderson et al., 2020) to approximate them:
$\frac{\partial u}{\partial x_{i}}({\bm{x}})\simeq\frac{u({\bm{x}}+h{\bm{e}}_{i})-u({\bm{x}})}{h},$ (22)
These approximations introduce errors, but the following result ensures that with a small enough $h$, these derivatives are very similar to the true ones.
Theorem C.1.
where ${\bm{e}}_{i}$ is the unit vector in the $x_{i}$ direction. Then, we have that $\|\mathrm{D}^{\epsilon}_{x_{i}}u\|_{L^{2}(\mathbb{R}^{n})}\leq\|\frac{\partial u}{\partial x_{i}}\|_{L^{2}(\mathbb{R}^{n})}$ and $\mathrm{D}^{\epsilon}_{x_{i}}u\rightarrow u_{x_{i}}$ in $L^{2}(\mathbb{R}^{n})$, which means that empirical derivatives converge to weak (or true) ones as $\epsilon\rightarrow 0$. The same results holds for any open set $\Omega$, with the convergence being true on every compact subset of $\Omega$. In practical terms, empirical derivatives converge a.e. for every point distant at least $h$ from the boundary.
@@ Appendix C On the use of Empirical Derivatives
Appendix D Additional Results and Experiments for Section 4

### v3

@@ 4.2 Generalization to new ODE/PDE parameters
We consider the dynamical system of a damped pendulum with state equation (DP) in table 1. Additional details on the experimental setup are available in Section C.4. We sampled a total of $50$ trajectories from different starting conditions $(u_{0},\dot{u}_{0})$, which act as parameters: $30$ are reserved for training, $10$ for validation, and $10$ for testing. Each trajectory was sampled every $\Delta t=0.01s$. The MLPs, which predict $u(t;u_{0},\dot{u}_{0})$ from $(t,u_{0},\dot{u}_{0})$, have 4 MLP with $50$ units each and $\tanh$ activation, and are trained for $200$ epochs with batch size $64$. Initial states are sampled randomly in $(u_{0},\dot{u}_{0})\in[-\frac{\pi}{2},\frac{\pi}{2}]\times[-1.5,1.5]$. In addition to the previously-used metrics and the angular velocities $\|\dot{u}-\hat{\dot{u}}\|_{2}$, we also monitor:
$\circ$ $\|\mathcal{G}[\hat{u}]\|_{2}$ (formally defined in Appendix C.4.1), the PDE residual which describes continuity and differentiability of trajectories with respect to the initial conditions. High values of $\|\mathcal{G}[\hat{u}]\|_{2}$ imply predicted trajectories do not change smoothly when varying $(u_{0},\dot{u}_{0})$, which is not consistent with the real system.
@@
For this task, we adapted Lagrangian Neural Networks (LNN) (Cranmer et al., 2019) and Hamiltonian Neural Networks (HNN) (Greydanus et al., 2019) to the damped pendulum case. We train them on the conservative part of the field (see C.4.2 for details), where they excel as they are specifically designed for conservative fields. Unlike DERL, they also require trajectories to be calculated through external solvers. Remarkably, DERL outperforms both LNN and HNN (figure 4(b)) on the learned field: DERL scores an error of $0.4445$, against $0.4470$ and $0.4496$ for HNN and LNN, respectively.
@@
Inspired by Penwarden et al. (2023), we consider a parametric experiment on the Allen-Cahn equation (AC). The parametric component is the external forcing $f$, calculated from the analytical solution $u(x,y;\xi)=e^{-\xi(x+0.7)}\sin(\pi x)\sin(\pi y)$, with parameter $\xi\in\mathbb{R}$. We reserved $8$ values of $\xi$ for training, $2$ for validation, and $5$ for testing, and we used $1000$ collocation points in the domain. The architecture is the same as in the Allen-Cahn experiment in Section 4.1. For further details and the complete setup, see Appendix C.5. Numerical results on the testing parameters are reported in Table 5, while further results are available in Appendix C.5.1. DERL always scores the best results, highlighting strong generalization abilities to unseen parameterization of the PDE. OUTL underperforms in both metrics, showcasing that learning from the raw data $u$ often leads to misleading solutions that do not follow the underlying physical laws. SOB underperforms as it again struggles to optimize a loss with both $u$ and $\mathrm{D}u$.
Double-parameter Allen-Cahn equation.
We now increase complexity by considering a dual-parameter version of the Allen-Cahn equation (Zou and Karniadakis, 2023). The analytical solution is given by $u(x,y;\xi_{1},\xi_{2})=\sum_{j=1}^{2}\xi_{j}\frac{\sin(j\pi x)\sin(j\pi y)}{j^{2}}$, from which the external forcing $f$ is calculated. We selected $15$ random couples of the parameters $(\xi_{1},\xi_{2})\in[0,1]^{2}$ for training, $5$ for validation and $5$ for test. Numerical results for the test parameters are reported in Table 5. Once again, DERL is consistently the best model for every couple of parameters (see Appendix C.5.2), performing at least two times better in predicting $u$ for unseen ones. Learning $u$ through OUTL leads to very poor performance in terms of physical consistency, even with the additional help of the PINN loss.
@@ 5 Transferring physical knowledge
We have shown that DERL is capable of extracting the relevant physical features from data.We now show how we can take advantage of this to transfer physical information between models through distillation (Hinton et al., 2015), eventually building physical models incrementally.
We set up our transfer experiments by dividing the whole training experience into 2 or 3 incremental steps. Each step extends the physical problem to i) new portions of the time domain, ii) to new portions of the space domain, and iii) to new parameterizations of the PDE. For these incremental training experiments, we assume an unsupervised setting where ground truth data is not available, and the models have to find a solution only by the definition of the PDE problem, as PINNs do.
@@
As a comparison, we train three additional PINN models. PINN-no-distillation updates the teacher with the PINN loss on the new domain, but without distillation on the previous domain, and is expected to forget previous information. PINN-full is a PINN trained on the full domain $\Omega$ without the iterative distillation process. This model is trained for the total sum of epochs of the other methodologies for a fair comparison. In continual learning, this joint training method is often taken as an upper bound for the expected performance of the final continual learning model Masana et al. (2023). Finally, we consider PINN-replay, which updates the teacher with the PINN loss on the new domain and on a subset of points (20%) from the previous domain. This is our closest resemblance to an Experience Replay setting.
@@ 5.1 Transfer to wider time horizons
In this first experiment with transferring, we aim to extend the working domain of a PINN in time. We consider the 1D Korteweg-de Vries equation (KdV) in Table 1. The spatial domain is the interval [1,1][-1,1]. while the entire time domain is $[0,1]$. We first train a PINN in the time interval $[0,0.35]$ and we collect its evaluations $\hat{u},\mathrm{D}\hat{u}$ on the training collocation points in a dataset. Then, a second model is trained with the combination of the PINN loss on the second time interval $[0.35,0.7]$, and a distillation loss on the old collocation points. This procedure aims to transfer old knowledge to this new model, while discovering the solution in a new interval of the time domain. Hence, this model should be able to learn the solution in the time-space domain $(t,x)\in[0,0.7]\times[-1,1]$. Finally, we repeat this procedure one more time, extending the model to the time interval $[0.7,1]$. As we will see, the final model has correctly learned the solution in the entire time-space domain of $0,1]\times[-1,1]$. For this task, each model is an MLP with $9$ layers of $50$ units and $\tanh$ activation, due to the more complex nature of the KdV task. Training phases last $30.000,30.000$, and $40.000$ steps each with the Adam optimizer and exponentially decaying lr, starting from $0.001$ with factor $0.9$. The PINN-full, PINN-no-distillation, and PINN-replay models train for the sum of these steps, that is $100.000$, with the same optimizer and learning rate schedule.
@@
PINN replay 5.371 1.889 5.281 10.54 13.72 6.289
Numerical results are available in Table 6. DERL outperforms all other methodologies in learning the solution to the PDE, either starting from-scratch or by fine-tuning the teacher model. Most importantly, our incremental methodology, coupled with the DERL distillation loss, improves the results by three times compared to a standard PINN. Figure 5 shows the prediction error in the whole domain as a function of the cumulative training steps for the continual models. A similar plot including the from-scratch models is available in Appendix D.3. These losses follow the typical behavior of continual learning, where the loss increases at the beginning of each phase and subsequently decreases, with the difference that performance is measured on the entire final task. The results clearly show that the PINN reaches a minimum at the beginning of training and cannot improve, while our incremental approach performs almost an order of magnitude better in the end. While PINN-replay improves on the baseline, using DERL to distill the teacher derivatives performs better to retain knowledge compared to experience replay. We notice that, even for transferring, it is not enough to distill the model’s outputs, both in terms of consistency and prediction error. Derivatives are both necessary and sufficient to distill the physical knowledge correctly. Finally, we see that incrementally training a PINN without distilling previous knowledge leads to forgetting of the learned solution. See Appendix D.3 for additional figures.
Figure 5: Prediction errors for the KdV transfer experiment for the continual models and the PINN variants. Vertical lines represent when a new step of incremental training is initiated.
5.2 Transfer to new domains
We leverage our transfer protocol to extend the working domain of a model (Figure 1). We consider the Allen-Cahn equation (AC) from Section 4.1. We first train a PINN in the lower-left quadrant of the domain ($1,0]\times[-1,0]$). We evaluate $\hat{u},\mathrm{D}\hat{u}$ on the training collocation points. Then, a second model is trained with a combination of the PINN loss on the lower-right quadrant ($0,1]\times[-1,0]$), and a distillation loss for the old points. The resulting model should be able to learn the solution in the lower half of the domain ($1,0]\times[1,0]$). The same procedure is repeated to extend the working domain of the model to the upper half ($1,1]\times[0,1]$). The final model should be a good predictor in the entire domain ($1,1]\times[-1,1]$). We used the same architecture as in Section 4.1 and $100$ epochs per step with the BFGS optimizer (Appendix D.4).
@@ 5.3 Transfer to new PDE parameters
Table 6 reports the $L^{2}$ errors and PDE residuals on the test parameters for both the from-scratch and continual cases. DERL proves to be the best distillation protocol. It also surpasses all PINN models, except for the full and replay PINNs on the PDE residual metric, whilst being the closest to match its performance. This again shows our method’s potential to transfer physical information across different instances of PDEs.
@@ 5.4 Distillation of higher-order derivatives
Finally, we investigate the effects of distilling higher-order derivatives. We consider the Korteweg-de Vries (KdV) equation (KdV in table 1), a third-order non-linear PDE. A pre-trained PINN teacher distills knowledge to students with identical architecture. We tested Hessian learning (HESL), which distills second-order derivatives and its combination with DERL (DERL+HESL) and SOB (SOB+HESL). The full setup and results can be found in section D.3. Here, we summarize the main findings. We observe that OUTL fails to distill effectively, as its PDE residual of $17.366$ is two orders of magnitude larger than the other models. Interestingly, the best-performing methods, HESL and DERL+HESL, did not access $u$ directly. They achieve the lowest PDE residual ($0.1915$ and $0.1931$, respectively) which is comparable to the PINN PDE residual of $0.1664$. Both approaches score the same error on $u$ as the teacher model, and even a lower BC loss.
@@ 6 Conclusion
We proposed DERL, a supervised methodology to learn physical systems from their partial derivatives. We showed theoretically and experimentally that our method successfully learns the solution to a problem and remains consistent with its physical constraints, outperforming other supervised learning methods. We also found DERL to be effective in transferring physical knowledge across models through distillation. We leveraged this capability to extend the working domain of a PINN and to generalize better across different PDE parameterizations. Finally, we showed the advantage of distilling higher-order derivatives. Our work shows promising intersections with continual learning, as DERL’s distillation allowed us to build physical models incrementally, showing minimal to no forgetting of previous knowledge. In the future, learning physical and dynamical systems may not always require starting from scratch. Rather, the learning process may be based on the continual composition and integration of physical information across different expert models. A more general and flexible paradigm that could lead to data-efficient and sustainable deep learning solutions.
Acknowledgments
This work has been supported by EU-EIC EMERGE (Grant No. 101070918).
@@ Appendix A Implementation Details and Setups
In theory, the augmented boundary should contain the values of $u(T,{\bm{x}})$ at the final time $T$. However, PDE problems are usually defined without a final condition, and it is more interesting not to have one to see which methodology can propagate physical information better. For example, SOTA methods as PINN do not include them (Raissi et al., 2019). Therefore, we decided not to include them as well, leading to a harder experimental setup than our theoretical results. For example, the experimental results in Appendix C.2 where the PINN failed to propagate the solution from $t=0$ to higher times, confirm the findings of Wang et al. (2022a) of PINNs being biased during training. However, DERL did not suffer from this issue.
@@ A.1 Computational time
Table 7: Computational time of 11 epoch for each experiment with the Adam optimizer. Allen-Cahn results are for the experimental setup in Appendix C.1. For NCL we reported the computational time of $100$ steps due to the definition and length of one epoch as in Richter-Powell et al. (2022).
@@ A.2 Model tuning
For the pendulum experiments that involve full trajectories, $30$ trajectories are for training and $10$ for validation, while $10$ unseen trajectories are used for model testing. For the pendulum interpolation task in Section C.4, a total of $40$ trajectories are used for in the tuning phase: points with times between $t=3.75$ and $t=6.25$ are unseen by the models and used for tuning/validation, while $10$ unseen complete trajectories are used for testing. For the parametric Allen-Cahn experiment with a single parameter, we calculated the solution on a set of $1000$ collocation points for $13$ values of $\xi\in[0,\pi]$: $8$ are used for training, $2$ for validation and $5$ for testing. Similarly, for the experiment with two parameters, we sampled $25$ couples of $(\xi_{1},\xi_{2})\in[0,1]^{2}$ with a $60/20/20$ split.
@@ Appendix B Proofs of the Theoretical Statements
We show the results for time-independent PDEs; for time-dependent PDEs, it is sufficient to consider the extended domain $\tilde{\Omega}=[0,T]\times\Omega$, in which the initial (and final) conditions become boundary conditions. We also show the result for functions with one component. The result can be generalized by applying it to each component.
Let $\hat{u}$ be the trained Neural network. From theorem 3.2, we have that a network $\hat{u}$ trained with the loss in equation 6 with $m$ derivative terms converges to the solution of the PDE $u$ in $W^{m,2}(\Omega)$. This implies that, when $L(\hat{u},u)\rightarrow 0$, all the derivatives of $\hat{u}$ converge to the true ones $D^{l}\hat{u}\rightarrow D^{l}u$ for $l=0,\ldots,m$. Since $\mathcal{F}$ is continuous, it follows that $\mathcal{F}[\hat{u}]\rightarrow\mathcal{F}[u]=0$. As a consequence, DERL learns to fulfill the underlying PDE of the system. ∎
@@
In the case of time-dependent PDEs, we will not provide the final conditions at $t=T$, which are part of the "boundary" of the extended domain and are necessary for Theorem 3.3. Since our experiments are on regular functions, this has proven not to be an issue, and the neural network still converges.
B.4 Proof of theorem 3.6
@@
Appendix C Additional Results and Experiments for Section 4